OpenStudy (anonymous):

Solve using the addition method. What is the x-value of the solution to the system? 2x – 3y = -1 3x + 4y = 24 4 3 2 none of the above

OpenStudy (anonymous):

2x-3y=-1_3x+4y=24 Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 12. 4*(2x-3y=-1)_3*(3x+4y=24) Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 12. 4*(2x-3y)=4(-1)_3*(3x+4y)=3(24) Multiply 4 by each term inside the parentheses. 4*(2x-3y)=-4_3*(3x+4y)=3(24) Multiply 4 by each term inside the parentheses. 8x-12y=-4_3*(3x+4y)=3(24) Multiply 3 by each term inside the parentheses. 8x-12y=-4_3*(3x+4y)=72 Multiply 3 by each term inside the parentheses. 8x-12y=-4_9x+12y=72 Add the two equations together to eliminate y from the system. 9x+12y=72_<U>8x-12y=-4<u>_17x =68 Divide each term in the equation by 17. x=4 Substitute the value found for x into the original equation to solve for y. 8(4)-12y=-4 Multiply 8 by each term inside the parentheses. 32-12y=-4 Move all terms not containing y to the right-hand side of the equation. -12y=-36 Divide each term in the equation by -12. y=3 This is the final solution to the independent system of equations. x=4_y=3